# Estimating the Influence of Sequentially Correlated Literary Properties in Textual Classification: A Data-Centric Hypothesis-Testing Approach

### v2

@@ Abstract
Stylometry aims to distinguish authors by analyzing literary traits assumed to reflect semi-conscious choices distinct from elements like genre or theme. However, these components often overlap, complicating text classification based solely on feature distributions. While some literary properties, such as thematic content, are likely to manifest as correlations between adjacent text units, others, like authorial style, may be independent thereof. We introduce a hypothesis-testing approach to evaluate the influence of sequentially correlated literary properties on text classification, aiming to determine when these correlations drive classification. Using a multivariate binary distribution, our method models sequential correlations between text units as a stochastic process, assessing the likelihood of clustering across varying adjacency scales. This enables us to examine whether classification is dominated by sequentially correlated properties or remains independent. In experiments on a diverse English prose corpus, our analysis integrates traditional and neural embeddings within supervised and unsupervised frameworks. Results demonstrate that our approach effectively identifies when textual classification is not primarily influenced by sequentially correlated literary properties, particularly in cases where texts differ in authorial style or genre rather than by a single author within a similar genre.
@@ 1 Introduction
As with many statistical-learning tasks, approaches in stylometry are usually categorized into supervised and unsupervised settings (Holmes, 1994, 1998; Juola, 2006). In the supervised case, several distinct authors and their works are provided (for training), and the task is to attribute texts to one of the authors under consideration, a task known as authorship attribution. Within this context, two distinct assumptions can be made: (1) that the author of each unattributed text is necessarily found within the given set of authors (Koppel et al., 2009), and (2) that the composer of an unattributed text need not be among the authors under consideration (Koppel et al., 2011). Alternatively, it may be desirable, given a single known author and some prior texts thereof, to verify whether some unknown work was also composed by this author or, instead, by a different one – a task known as authorship verification (e.g., Koppel et al., 2007; Juola, 2006; Kestemont et al., 2016). Other supervised settings include authorship profiling, where the labeling required is not for a distinct author but for a group with some shared profile (e.g., historical era, native language, gender, socio-economic class) (Argamon et al., 2009). Furthermore, variants of these tasks include cross-genre settings, where the train set is from a different genre than the test set (e.g., prose-poetry (Kestemont et al., 2012)), cross-domain settings (e.g., Facebook-Twitter Markov et al. (2018)), or both (Stamatatos, 2018). Lastly, there exists a variety of different case studies that incorporate different languages, historical eras, text sizes, and more (Schwartz et al., 2013).
In the unsupervised setting, no prior knowledge of the authors is provided. For a corpus of unlabeled texts (or a single text split into text units, such as sentences, paragraphs, or some arbitrary partition), it is required to estimate which texts were composed by the same author. At the root of this unsupervised task is the fundamental question: Given two texts, did the same author write them or not?
As the challenge at hand is identifying a typical style regardless of thematic content, literary features used in stylometry focus on the abundant rather than the scarce (Koppel and Winter, 2014). Unique standing-out words that allude to content are usually ignored, and preferred features are traditionally subsets of the most frequent words and sequences thereof, as these could attest to some authorial style that persists through different works regardless of content (e.g., Holmes, 1994; Argamon et al., 2007; Stamatatos, 2009; Escalante et al., 2011). Even today, when an interpretable analysis is of importance, commonly used features are $n$-gram features of words, characters ($k$-mers), part-of-speech tags (POS), or dependency parsing tags, which are usually applied with a frequency-dependent vectorization such as tf-idf (Sammut and Webb, 2010).
While some attention has been devoted to a thorough assessment of the utility of different literary features in various settings, most of this attention has focused on variants of the supervised setting of stylometry (Koppel et al., 2013; Eder, 2015; Juola et al., 2019). In the unsupervised setting, much less has been studied about the effects of different linguistic features and their dependence on textual properties such as genre or language-specificity. Moreover, the focus has been on assessing the contribution of different features and methods in achieving benchmarks. However, to the best of our knowledge, no prior work deals with ascertaining what part of the signal found can verifiably be attributed to different constituents of the literary signal, which can lead to false assertions concerning the results of such analyses.
Ensuring that an authorship attribution algorithm remains agnostic of qualities of the text that may be independent of style, such as theme or genre, is crucial for robust and accurate analysis (e.g., Eder, 2015; Rudman, 2016; Murauer and Specht, 2021). Thus, the algorithm can focus solely on identifying each author’s unique stylistic and linguistic patterns. This approach guards against biases arising from thematic similarities or differences between texts, which could obscure genuine authorial signatures. However, achieving such agnosticism presents challenges. Texts encompass a multitude of features, and separating the relevant stylistic cues from the thematic or genre-related elements – that are often closely coupled to style, requires sophisticated algorithms capable of discerning subtle linguistic nuances (e.g., Halvani et al., 2020). Moreover, a supervised algorithm must be trained on diverse datasets spanning various themes and genres to ensure its adaptability and effectiveness across different textual domains. Furthermore, employing interpretable methods is essential to ensure that the algorithm performs according to expectations, allowing researchers to understand and validate the reasoning behind its attribution decisions (e.g., Theophilo et al., 2022). Thus, balancing the need for agnosticism of non-stylistic content with the complexity of textual data poses a significant obstacle in developing authorship attribution algorithms that are both reliable and versatile.
In this work, we present and explore a novel approach that takes a first step towards identifying which constituent(s) of the literary signal (e.g., thematic, stylometric, genre-wise) is responsible for a distinction between a pair of texts. Specifically, we wish to rule out that the distinction is based on constituents that exhibit correlations between adjacent text units that include but are not limited to thematic content. Our method employs a hypothesis-testing framework that analyzes sequential correlations of label sequences of text units. By developing a stochastic model for these correlations, we provide an empirical basis for assessing whether sequentially correlated literary properties dominate classification outcomes. We demonstrate that relying solely on embeddings intended to capture stylistic features is often insufficient to rule out non-stylistic influences, especially when text units are short. Instead, our approach models sequential correlation distributions independently of any specific embedding or algorithm. Since adjacent text units tend to be correlated by, for example, thematic content, rather than necessarily by authorial style (in the event of distinct authors), they naturally exhibit similarity when embedded.
We consider pairs of texts made of text units of English prose of different similarity in genre, composed by either other or the same author. We conduct a lateral study by applying unsupervised and golden-standard supervised methods to distinguish between pairs of texts embedded using traditional and state-of-the-art neural embeddings optimized for authorship attribution. In each case, we apply our hypothesis-testing framework to determine whether the resulting classification is predominantly affected by sequential correlations – as we would expect for two texts written by the same author and are similar in genre and are therefore likely to cluster based on thematic distinctions – or not.
@@ 2 The State of Stylometry
Stylometry can be traced to the mid-20th century, when scholars such as Mosteller and Wallace employed statistical methods to analyze the writing styles of the Federalist Papers, demonstrating the effectiveness of word frequency and function word usage in distinguishing between authors (Mosteller and Wallace, 1963). Their pioneering work set the stage for applying machine learning and natural language processing techniques in stylometric research.
Advancements in computational linguistics and data science in the late 20th and early 21st centuries increased the scope of stylometric analysis expanded beyond authorship attribution to include authorship verification, genre classification, and text clustering (e.g., Juola et al., 2006; Juola, 2006; Rybicki and Heydel, 2013).
One of the landmark studies in contemporary stylometry by Koppel et al. (2007) explored the stylistic fingerprints of authors across different languages and genres. Using machine learning algorithms and linguistic features, they demonstrated the robustness of stylometric techniques in identifying individual writing styles and distinguishing between authors with high accuracy. There, they demonstrated and emphasized the importance of frequent features, such as common words or phrases, punctuation patterns (Darmon et al., 2021), and syntactic structures, in capturing the unique writing style of an author. They argued that these frequent features are more indicative of an author’s stylistic tendencies than rare or unique words, which may be more subject to variation or noise. Later works underpinned the fact that frequent features alone may not capture style comprehensively (e.g., Savoy, 2013; Lagutina et al., 2019; Grieve, 2023), and highlighted the importance of considering the limitations of frequency-based stylometric features and the need for complementary approaches to capture the richness and complexity of writing style.
Complementary methods to capture writing style beyond frequent features have been proposed to address the limitations of frequency-based stylometric analysis. One such approach involves incorporating syntactic and structural features of text, which can provide deeper insights into an author’s unique stylistic tendencies. Further studies have demonstrated the effectiveness of syntactic features, such as parse tree structures (e.g., Feng et al., 2012) and syntactic dependencies (e.g., Hollingsworth, 2012), in distinguishing between authors and genres. Additionally, semantic analysis techniques, which focus on the meaning and context of words and phrases, have been proposed as valuable complements to frequency-based methods. In conclusion, by integrating these complementary methods, researchers can achieve more comprehensive and accurate analyses of writing style in various textual domains.
More recently, stylometry experienced a significant evolution with the advent of Large Language Models (LLMs) and neural approaches (e.g., Ding et al., 2017; Canbay et al., 2020). Leveraging the power of large-scale pre-trained models such as the Generative Pre-trained Transformer (GPT) series, researchers have explored novel avenues in authorship attribution, genre classification, and other stylometric tasks (e.g., Huertas-Tato et al., 2024). Recent works showcase the capabilities of LLMs when applied to the authorship attribution task (e.g., Kumarage and Liu, 2023). However, despite their impressive performance in a wide variety of disciplines, as well as within the domain of language processing, many of these neural approaches lack interpretability, making it challenging to understand the underlying factors driving their predictions (Dror et al., 2018, 2020; Moulaison-Sandy, 2023). Moreover, in contexts such as historical texts where training data may be scarce or non-representative, these sophisticated models may not be applicable, highlighting the need for interpretable and domain-adaptive stylometric methods.
Despite the rapid and impressive growth in textual analysis for a myriad of purposes, there has been a noticeable scarcity of works specifically aimed at discerning various literary parameters, including theme, style, and genre, to ascertain which one predominantly drives text classification (Lagutina et al., 2019; Litvinova, 2020; Hou and Huang, 2020). This lack of focus on disentangling the distinct influences of these parameters – which is difficult and often impossible due to the inherent vaguely defined characteristics thereof – can lead to potential false positive classifications based solely on a hypothesized factor (e.g., Schuster et al., 2020). Therefore, while stylometric analysis traditionally prioritized tasks such as authorship attribution, a broader consideration of literary parameters is essential for accurate text classification and a deeper understanding of textual characteristics.
@@ 3 Hypothesis-Testing Scheme
In the context of textual data analysis, sequential correlations are believed to manifest at different adjacency scales due to the inherent structural organization of the text. Textual content often exhibits sequential coherence and organization, where related concepts, ideas, or themes are often clustered together within contiguous segments of text (e.g., Ghadessy and Gao, 2001; Kim and Wilbur, 2012; Armborst, 2017). At a finer adjacency scale, adjacent words or phrases within a sentence or paragraph are more likely to be sequentially related, capturing local textual coherence. On larger adjacency scales, such as between sentences, paragraphs, or chapters, thematic correlations may evolve to capture broader sequential shifts or transitions within the text. For instance, transitions between paragraphs may signal shifts in topic or theme, while transitions between chapters may denote major shifts in narrative focus or subject matter. By considering sequential correlations at these different adjacency scales, we can gain a more comprehensive understanding of the text’s underlying sequential structure and organization, enabling more accurate analysis and interpretation of its classification.
Recently, we suggested in Yoffe et al. (2023) that correlations between adjacent text units affect their classification based on thematic similarity, potentially masking stylistic differences. Consequently, when devising a hypothesis-testing framework to estimate the statistical significance of a potentially stylistic distinction between texts, the assumption of naive label permutation, which ignores these correlations and assumes label independence, is unsuitable for textual data. We proposed that human experts’ labeling tendencies may reflect a predisposition to assign authorship to text units sharing sequential similarities, as often observed in Biblical exegetical analysis (e.g., Dershowitz et al., 2015). This suggests that when human experts annotate text segments, they tend to cluster together words, phrases, or passages with thematic coherence, interpreting them as originating from the same authorial source. There, we made the statistical argument that shifting the labels cyclically conserves predominantly sequential correlations. Indeed, this cyclic-shift test is a parochial solution and is often unsuitable for similar tasks since even if a human expert’s labeling is available – its underlying rationale can be affected by a convoluted mixture of literary parameters.
Here, we propose a data-centric approach that assumes labels to be correlated random variables in the sequential domain (what we refer to as label space). By only being concerned with label sequences, this approach is essentially embedding-invariant and can be applied to texts (or other datasets) embedded densely, sparsely, or with any desired subset of features or neural or traditional language model. By estimating the intrinsic average correlations between text units at different lags (i.e., distances along the textual sequence), we generate a multivariate distribution from which we draw label sequences that are accordingly correlated between text units at different sequential scales. This approach enables us to sample a random set of labels in a way that simulates the inherent sequential correlations apparent in texts and, therefore, provides a null hypothesis with such characteristics.
In this work, we consider the null hypothesis labeling-generation routine for the simplest scenario of a text composed by one or two authors. Once this sampling mechanism for our null distribution is established, we can compute the $p$-value indicating the likelihood that the distinction between the two parts of the text indicated by the labels of $L$ is made due to classification driven by sequentially correlated literary properties.
Reference label sequence $L^{(\rm{real})}$ of a textual data set
We consider correlations in the label space. Let $D\in\mathbb{R}^{m\times f}$ denote a textual data set containing $m$ samples (text units – a sequence of words of some length) of $f$ dimensions (features). We denote a sequence of labels representing some hypothesized or real partition between texts (or an internal partition within a single text, as discussed in Yoffe et al. (2023)) as $L^{(\rm{real})}$.
Label sequence $L$ of a classification algorithm applied to $D$
@@
Average auto-covariance matrix $M$
@@
Intuition: The average autocovariance matrix $M$ serves as a structured representation of the typical correlations (i.e., $A_{\ell}$) between labels at varying sequential distances ($\ell$) within the dataset. The $i$th row in $M$ corresponds to the linearly-decaying autocovariance of the $i$th label ($(M)_{ii}$) with respect to all other labels in the dataset ($(M)_{ij\neq i}$), which is governed by (1): their proximity to it (i.e., lag), and (2) the average autocovariance between all labels in the dataset separated by that lag. This means that entries closer to $(M)_{ii}$ represent correlations between the $i$th label and its closer neighbors, and they would therefore be more correlated thereto than farther entries. Thus, $M$ specifically captures the average autocovariance between labels at different sequential distances within the dataset – the necessary mechanism to model sequential correlations. As opposed to $M$, a simple covariance matrix captures the covariance between pairs of labels in the dataset without considering their sequential arrangement. Each entry in the covariance matrix quantifies pairwise relationships between every two labels, irrespective of their positions within the sequence.
Generating a correlated label sequence $L^{(null)}$
@@
We use the autocovariance matrix $M$ to draw a random sequence of correlated labels $L^{(\rm{null})}\in\mathbb{R}^{m}$ in the following manner: consider the multivariate normal distribution $V\sim\mathcal{N}(\overrightarrow{\bar{L}},M)$, where $\overrightarrow{\bar{L}}=(\bar{L},\ldots,\bar{L})^{T}\in\mathbb{R}^{m}$. The choice of $V$ as a multivariate normal distribution stems from the ease of embedding the average autocovariance matrix $M$ as a covariance matrix. Let $X=(X_{1},...,X_{m})\in\mathbb{R}^{m}$ be drawn from $V$, which can be transformed into a vector of correlated labels ($L^{(\rm{null})}$) using elementwise thresholding (Leisch et al., 1998):
$L_{i}^{(\rm{null})}=\begin{cases}1,&X_{i}>0.5\\ 0,&\textrm{otherwise}.\end{cases}$
@@
Essentially, this means that the probability $p_{i}$ of the $i$th label in $L^{(\rm{null})}$ to be 1 is given by
$p_{i}\equiv\mathbb{P}(L^{(\rm{null})}_{i}=1)=\mathbb{P}(X_{i}>0.5).$ (3)
Intuition
The multivariate normal distribution $V$ is a convenient method to couple the two elements that are essential for generating stochastic correlated label sequences: (1) the probability of each label to be 1, which is identical for all entries and is simply the average label value of the reference label sequence $L$, and (2) the average autocovariance matrix $M$, determining how correlated each entry in the vector $X$ that is drawn from $V$ is with all other labels. $X$, in turn, is a specific realization drawn from $V$, whose each entry is a continuous number, affected by correlations with the entire dataset. Lastly, elementwise thresholding ensures that the values of $X$ assume the binary form of a label sequence – either 0 or 1.
Thus, $V$ enables the generation of stochastic label sequences that preserve sequential dependencies that exist in the reference label sequence $L$, thereby facilitating a robust method to generate a null distribution that would examine the measure of agreement between sequentially correlated label sequences and a real or hypothesized partition between two texts.
Within a hypothesis-testing framework, we may probe the null hypothesis whereby the reference label sequence $L$ is one such realization of labels that only exhibit sequential correlations. Thus, assuming that $L$ is characterized primarily by sequential correlations, we use it to generate the underlying distribution from which it was presumably sampled. Using this distribution, we generate stochastic label sequences that possess similar sequential qualities. In the event that the agreement between the real label sequence $L^{(\rm{real})}$ reference label sequence $L$ is not based on sequential correlations (but rather on, e.g., stylistic differences) – that is, that the null hypothesis is wrong, the resulting overlap between the null label sequences and $L^{(\rm{real})}$ will not be comparable to its overlap with $L$ since the bulk of agreement between the two originates from a non-sequential axis. This comparison is insightful regarding the extent to which sequential correlations contribute to the observed label assignments, underpinning the robustness and significance of non-sequentially correlated distinctions between texts. In Figure 1, we provide a visualization of $A$ and $M$ for some label sequence $L$ of length 1300 to provide some additional intuition.
Figure 1: Vizualization of the hypothesis-testing framework’s building-blocks: Left Panel: Average auto-covariance vector $A$ computed for some label sequence of size 1300. The $i$th entry is $A_{\ell=i}$ (see Eq. 1). Right Panel: Average auto-covariance matrix $M$ (see Eq. 2) The insert visualizes the 500th row in $M$, which is a mirror image of $A$ around the 500th entry, mapping correlations between that label and all other labels in the sequence.
Matthews Correlation Coefficient (MCC)

### v4

@@ Abstract
We introduce a data-centric hypothesis-testing framework to quantify the influence of sequentially correlated literary properties—such as thematic continuity—on textual classification tasks. Our method models label sequences as stochastic processes and uses an empirical autocovariance matrix to generate surrogate labelings that preserve sequential dependencies. This enables statistical testing to determine whether classification outcomes are primarily driven by thematic structure or by non-sequential features like authorial style.
Applying this framework across a diverse corpus of English prose, we compare traditional (word $n$-grams and character $k$-mers) and neural (contrastively trained) embeddings in both supervised and unsupervised classification settings. Crucially, our method identifies when classifications are confounded by sequentially correlated similarity, revealing that supervised and neural models are more prone to false positives – mistaking shared themes and cross-genre differences for stylistic signals. In contrast, unsupervised models using traditional features often yield high true positive rates with minimal false positives, especially in genre-consistent settings.
By disentangling sequential from non-sequential influences, our approach provides a principled way to assess and interpret classification reliability. This is particularly impactful for authorship attribution, forensic linguistics, and the analysis of redacted or composite texts, where conventional methods may conflate theme with style. Our results demonstrate that controlling for sequential correlation is essential for reducing false positives and ensuring that classification outcomes reflect genuine stylistic distinctions.
@@ 1 Introduction
Stylometric approaches are typically categorized into supervised and unsupervised settings (Juola et al., 2008). In supervised tasks, a set of authors and their works are provided for training, with the goal of attributing an unattributed text to one of the authors—known as authorship attribution. Two assumptions can apply: (1) the author must be among the provided set (Koppel et al., 2009), or (2) the author may not be in the set (Koppel et al., 2011). Alternatively, authorship verification aims to determine if an unknown work was written by a known author or someone else (e.g., Koppel et al., 2007; Juola et al., 2008; Kestemont et al., 2016). Other supervised tasks include authorship profiling, where texts are attributed to a shared group profile (e.g., historical era, gender, language) (Argamon et al., 2009), as well as cross-genre and cross-domain settings (e.g., Kestemont et al., 2012; Stamatatos, 2018). These tasks have been applied across different languages, eras, and text sizes (Schwartz et al., 2013).
In the unsupervised setting, no prior author information is given. The task is to identify whether pairs of unlabeled texts (or text units like sentences or paragraphs) were written by the same author. The core question is: Did the same author write these texts?
This work introduces a novel approach to identifying whether sequentially correlated literary properties contribute to text classification. We apply a hypothesis-testing framework to analyze such correlations in label sequences, which are often linked to thematic content, rather than authorial style. Our stochastic model of sequential correlations offers an empirical basis for distinguishing between stylistic versus certain non-stylistic influences.
We test our approach on English prose texts of varying genre similarity, authored either by the same or by different authors. Using both unsupervised and supervised methods with traditional and state-of-the-art neural embeddings – optimized for capturing authorial style, we apply our hypothesis-testing framework to assess whether the classification is primarily driven by sequential correlations, as seen in same-author, genre-similar texts, or by other factors. These experiments affirm that accounting for the influence of sequential correlations in textual classification consistently improves the distinction between stylistic and non-stylistic classification outcomes in all the setups we tried.
This research is motivated by the challenges posed by multilayered and redacted historical texts, which require interdisciplinary approaches across linguistics, history, and computer science to trace their evolution from origin to extant form (e.g., Bühler et al., 2024). These texts frequently include annotations, marginalia, and modifications made by scribes or editors, obscuring the original authorial intent (e.g., Müller et al., 2014), and are often written in archaic or obsolete languages that complicate automated analysis (e.g., Baum, 2017). Additionally, their interpretation demands careful consideration of historical context, including cultural norms, linguistic conventions, and socio-political factors, to ensure accurate analysis (Chapman, 2017). While our current work does not directly address historical texts, it seeks to validate our approach by applying it to modern texts with undisputed authorial and stylistic qualities, thereby providing a robust foundation for future applications to more complex historical material.
@@ 2 The State of Stylometry
Stylometry, the quantitative study of literary style, can be traced to the mid-20th century when scholars such as Mosteller and Wallace applied statistical methods to assess the authorship of the Federalist Papers, demonstrating the effectiveness of function word frequencies and statistical inference in distinguishing authorship (Mosteller and Wallace, 1963). This pioneering work provided early evidence that consistent, measurable linguistic signals could be exploited to resolve long-standing literary and historical debates, laying the groundwork for contemporary computational stylometry.
With the growth of computational linguistics and data-driven approaches in the late 20th and early 21st centuries, the scope of stylometric research broadened considerably. Beyond basic authorship attribution, researchers developed models for authorship verification, author profiling, genre classification, and clustering of anonymous or disputed texts (e.g., Juola et al., 2006, 2008; Rybicki and Heydel, 2013). One influential contribution from this period is Burrows’ Delta method (Burrows, 2002), which quantifies stylistic distance using Z-score-based comparisons of function word frequencies. Its simplicity, language independence, and empirical robustness have made it a foundational benchmark in both literary and forensic stylometry. As a widely adopted baseline for measuring authorial similarity, Delta offers a natural point of comparison for the present work, which introduces a sequential correlation-based framework aimed at disentangling stylistic signals from theme-driven textual structure.
A landmark machine learning-oriented approach by Koppel et al. (2007) demonstrated the robustness of stylometric classification using frequent features such as common lexical items, function words, and character-level statistics. Their work laid the foundation for high-accuracy authorship identification across genres and languages. Further refinements—incorporating punctuation patterns (Darmon et al., 2021) and syntactic features—showed that such frequent signals could capture authorial style more reliably than topic- or content-driven lexical features, which are often more variable and less discriminative.
Nonetheless, growing evidence has shown that traditional frequency-based features can be confounded by thematic content. Mikros and Argiri (2007) demonstrated that many widely-used stylometric variables are topic-sensitive, and may mislead attribution systems when stylistic and thematic patterns overlap. This has led to a critical reassessment of assumptions underlying feature selection. Studies such as Savoy (2013), Lagutina et al. (2019), and Grieve (2023) argue that true stylistic modeling requires more than frequency analysis—it requires incorporating structural, syntactic, and discourse-level features capable of representing deeper authorial patterns.
To address these limitations, several works have proposed syntactic and structural approaches. For example, Feng et al. (2012) employed syntactic parse tree structures to detect deception and authorial signals, while Hollingsworth (2012) showed that dependency parsing can help distinguish authorial style from topic. These syntactic methods provide a level of abstraction that is less sensitive to thematic variation and more robust across genres and text types.
Information-theoretic and compression-based approaches have also gained traction, particularly due to their robustness and language-agnostic nature. Cilibrasi and Vitányi (2005) introduced normalized compression distance (NCD), a compression-based measure of similarity that operates without predefined linguistic features or language-specific preprocessing. By quantifying the shared information content between two texts using standard compression algorithms, NCD provides a robust, domain-agnostic method for clustering and comparison. This approach offers a valuable alternative to traditional feature-engineered methods, and its emphasis on holistic textual similarity complements the data-centric perspective of the current work—particularly in highlighting classification signals that emerge from structural regularities rather than explicitly engineered feature sets.
Stylometry has also become increasingly relevant in forensic linguistics, where the need for interpretable, defensible results is paramount. Juola (2007) emphasized the importance of transparency and reproducibility in authorship analysis, especially in legal or historical contexts where the evidentiary burden is high. More recent work has focused on evaluating the reliability and statistical significance of classification results, often calling for hypothesis-driven models that can quantify uncertainty and isolate confounding factors.
The rise of neural language models has brought a new wave of high-performance tools to stylometry. Studies such as Ding et al. (2017) and Canbay et al. (2020) explore deep learning models trained for stylometric tasks, including authorship attribution and genre classification. These models leverage high-dimensional contextual embeddings, often derived from large pre-trained transformers. While they show strong empirical results, they are frequently criticized for their lack of interpretability and for their potential sensitivity to thematic and structural artifacts (Dror et al., 2018, 2020).
Despite these methodological advances, a major gap persists in understanding what type of information—style, theme, or genre—drives classification success. Litvinova (2020), Hou and Huang (2020), and Schuster et al. (2020) highlight the risks of misattributing thematic similarity as stylistic consistency. This concern is especially salient in historical, redacted, or multi-author texts where thematic coherence may coexist with significant stylistic variation. Without tools to differentiate these axes, stylometric analyses risk drawing misleading conclusions.
This distinction is especially crucial in the analysis of complex or redacted texts, such as those found in the biblical corpus. Many of the books in the Hebrew Bible, for instance, are widely believed to be composite, incorporating contributions from multiple authorial sources, traditions, and editorial strata over time (e.g., Wellhausen, 1885; Holzinger, 1893; Gunkel, 1895). Within the framework of the documentary hypothesis, texts such as the Book of Exodus and the Book of Genesis have been segmented into distinct sources (e.g., Priestly, Elohistic, Yahwistic) based on recurring stylistic, lexical, and thematic features. Yet, these divisions are often based on assumptions that thematic consistency necessarily reflects authorship or redaction (Wellhausen, 1885; Albertz, 2018). Recent work has sought to bring more computational rigor to these debates. For example, Dershowitz et al. (2015), Yoffe et al. (2023), Bühler et al. (2024), and Faigenbaum-Golovin et al. (2024) employed stylometric clustering to detect internal textual divisions corresponding to hypothesized strata—but even these results are subject to interpretative ambiguity, as stylistic boundaries may be blurred by thematic layering or editorial interpolation.
A related example comes from the domain of classical literature. In their stylometric analysis of Latin texts, Kestemont et al. (2016) developed an authorship verification framework to assess the authenticity of disputed works attributed to Julius Caesar. By training classifiers on known Caesar texts and testing against contested pieces, they achieved strong empirical results—but the interpretation of these classifications again hinged on whether the distinguishing signals reflected Caesar’s unique stylistic fingerprint or domain-specific content (e.g., military vocabulary, rhetorical form). While their framework operated under the assumption that thematic bias could be mitigated through feature design and normalization, our approach would enable a more explicit assessment: namely, whether the classification boundaries align with patterns of sequential thematic development rather than true stylistic divergence.
In both historical-critical and classical philological settings, the challenge remains the same: when classification succeeds, what exactly is being distinguished? Our hypothesis-testing framework is designed to answer this question directly. By modeling the expected distribution of label sequences that arise from sequentially structured (but non-stylistic) sources, we provide a principled way to separate theme-driven partitions from genuinely stylistic ones. This opens the door for more cautious and informed interpretations of stylometric findings, particularly in cases where authorial boundaries are unknown or the text has been shaped by centuries of redaction, translation, and reuse.
In response to this gap, the present work introduces a hypothesis-testing framework for assessing the influence of sequentially correlated properties—such as theme—on classification outcomes. By generating surrogate label sequences that preserve observed autocorrelation patterns, we estimate whether classification results are likely driven by sequential structure or by non-sequential stylistic signals. This approach complements existing stylometric methods by offering a data-centric, statistically grounded test of interpretability—particularly valuable in contexts where both style and theme may be entangled.
@@ 3 Hypothesis-Testing Scheme
We propose a data-centric approach that treats labels as correlated random variables in the sequential domain. This embedding-invariant method can be applied to texts (or other datasets) embedded densely, sparsely, or with any subset of features from neural or traditional language models. By estimating the intrinsic average correlations between text units at different lags (i.e., distances along the textual sequence), we generate a multivariate distribution from which we draw label sequences that reflect these sequential correlations. This provides a null hypothesis that simulates the inherent sequential dependencies in the text.
We consider the null hypothesis labeling-generation routine for the simplest scenario of a text composed by one or two authors. Once this sampling mechanism for our null distribution is established, we can compute the $p$-value indicating the likelihood that the distinction between the two parts of the text indicated by labels $L$ is made due to classification driven by sequentially correlated literary properties.
Reference label sequence $L^{(\mathrm{real})}$ of a textual data set.
We consider correlations in the label space. Let $D\in\mathbb{R}^{m\times f}$ denote a textual data set containing $m$ samples (text units – a sequence of words of some length) of $f$ dimensions (features). We denote a sequence of labels representing some hypothesized or real partition between or within texts as $L^{(\mathrm{real})}$.
Label sequence $L$ of a classification algorithm applied to $D$.
@@
Average auto-covariance matrix $M$.
@@
Generating a correlated label sequence $L^{(null)}$.
@@
We use the autocovariance matrix $M$ to draw a random sequence of correlated labels $L^{(\mathrm{null})}\in\mathbb{R}^{m}$ in the following manner: consider the multivariate normal distribution $V\sim\mathcal{N}(\overrightarrow{\bar{L}},M)$, where $\overrightarrow{\bar{L}}=(\bar{L},\ldots,\bar{L})^{T}\in\mathbb{R}^{m}$. The choice of $V$ as a multivariate normal distribution stems from the ease of embedding the average autocovariance matrix $M$ as a covariance matrix. Let $X=(X_{1},...,X_{m})\in\mathbb{R}^{m}$ be drawn from $V$, which can be transformed into a vector of correlated labels ($L^{(\mathrm{null})}$) using element-wise thresholding (Leisch et al., 1998):
$L_{i}^{(\mathrm{null})}=\begin{cases}1,&X_{i}>0.5\\ 0,&\textrm{otherwise}.\end{cases}$
@@
Essentially, this means that the probability $p_{i}$ of the $i$th label in $L^{(\mathrm{null})}$ to be 1 is given by
$p_{i}\equiv\mathbb{P}(L^{(\mathrm{null})}_{i}=1)=\mathbb{P}(X_{i}>0.5).$ (3)
Intuition.
The multivariate normal distribution $V$ offers a convenient and analytically tractable method for generating stochastic binary label sequences that preserve the sequential dependencies observed in a reference sequence $L$. This model couples two essential components: (1) the marginal probability of a label being 1, which is assumed to be constant across entries and estimated by the empirical mean of $L$ (i.e., $\bar{L}$), and (2) the average autocovariance matrix $M$, which encodes the expected correlation between each label and every other label in the sequence based solely on their relative distance.
The matrix $M$ is constructed as a symmetric Toeplitz matrix, with entries defined by $(M)_{ij}=A_{|i-j|}$, where $A$ is the vector of average autocovariance across all lags. This Toeplitz structure reflects an important modeling assumption: that the correlation between labels is translation-invariant—that is, the degree of correlation between any two labels depends only on how far apart they are, not on their absolute position in the sequence. This aligns with the hypothesis that thematic or stylistic coherence decays smoothly with distance in natural language, regardless of context or boundary location.
A sample vector $X$ drawn from $V$ is thus a continuous-valued realization that preserves the second-order (i.e., pairwise) sequential correlations of $L$. To convert this to a binary label sequence, we apply element-wise thresholding (e.g., assigning a value of 1 if $X_{i}>0.5$ and 0 otherwise). This produces a new binary sequence that shares the same correlation structure as the original, without inheriting any specific content or higher-order dependencies.
Within a hypothesis-testing framework, this allows us to define a null hypothesis: that the observed reference label sequence $L$ reflects nothing more than the natural sequential correlations found in the text. Under this null, $L$ is treated as a representative draw from the multivariate normal model $V$, which is then used to simulate a distribution of surrogate label sequences possessing similar sequential properties.
We then evaluate the degree of agreement between each surrogate sequence and a target label sequence $L^{(\mathrm{real})}$—which may represent a hypothesized or externally assigned partition of the text (e.g., by authorship or stylistic phase). If the agreement between $L$ and $L^{(\mathrm{real})}$ can be explained by sequential correlations alone, then the overlap between the surrogate sequences and $L^{(\mathrm{real})}$ should be statistically comparable. However, if the overlap between $L$ and $L^{(\mathrm{real})}$ significantly exceeds what is expected under the null, we infer that their agreement likely reflects a meaningful, non-sequential structure—such as a real authorial distinction.
This approach thus enables a principled test of whether the separability in a classification task is driven by underlying literary structure rather than being an artifact of natural sequential coherence in language. In Figure 1, we visualize the autocovariance vector $A$ and matrix $M$ computed from a label sequence $L$ of length 1213, illustrating the sequential correlation structure modeled in our hypothesis-testing framework.
Figure 1: Flowchart of the hypothesis-testing framework. Starting from an embedded text corpus $D\in\mathbb{R}^{m\times n}$, a binary label sequence $L\in\{0,1\}^{m}$ is obtained via 2-means clustering. The autocovariance vector $A$ and corresponding matrix $M$ are computed from $L$. A stochastic vector $V\sim\mathcal{N}(\vec{\bar{L}},M)$ is drawn and thresholded to generate null sequences $L^{(\text{null})}$. Repeating this process yields a null distribution of MCC scores against $L$, from which a $p$-value is estimated.
Matthews correlation coefficient (MCC).